# Where Did This Sentence Come From? Tracing Provenance in LLM Reasoning Distillation

### v1

@@ A.6.2 Model Auditing
Another closely related area is model auditing, a growing line of work that studies (model_audit1; model_audit2; model_audit3) auditing generative models to understand what data they memorize and to attribute outputs back to underlying data sources. For example, prior work (model_audit1) shows that rare tokens in the training data tend to be memorized by text generation models, and uses shadow models together with an audit classifier (e.g., an SVM on token-rank features) to distinguish whether a user’s data was included in training. Separately, subsequent work (model_audit2) formalizes extractability as the ability of a model, given a prefix, to greedily regenerate the exact suffix from the training set, and systematically studies how repetition and sequence length affect the fraction of such extractable sequences. In contrast, our work targets model-level provenance in a distillation setting: rather than asking whether specific data are memorized, we aim to trace which upstream models are the sources of a given output, shifting the focus from data membership to the lineage of the models themselves.

### v2

@@ A.6.2 Model Auditing
Another closely related area is model auditing, a growing line of work that studies (Song and Shmatikov, 2019; Carlini et al., 2022; Dekoninck et al., 2024) auditing generative models to understand what data they memorize and to attribute outputs back to underlying data sources. For example, prior work (Song and Shmatikov, 2019) shows that rare tokens in the training data tend to be memorized by text generation models, and uses shadow models together with an audit classifier (e.g., an SVM on token-rank features) to distinguish whether a user’s data was included in training. Separately, subsequent work (Carlini et al., 2022) formalizes extractability as the ability of a model, given a prefix, to greedily regenerate the exact suffix from the training set, and systematically studies how repetition and sequence length affect the fraction of such extractable sequences. In contrast, our work targets model-level provenance in a distillation setting: rather than asking whether specific data are memorized, we aim to trace which upstream models are the sources of a given output, shifting the focus from data membership to the lineage of the models themselves.